# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v3

@@ III-C1 Assumptions on ($\mathcal{P}$)
$\left\|\nabla f({\mathbf{x}},{\boldsymbol{\xi}})-\nabla f({\mathbf{y}},{\boldsymbol{\xi}})\right\|$ $\leq L\left\|{\mathbf{x}}-{\mathbf{y}}\right\|$ (12)
@@
$\lvert g_{j}({\mathbf{x}})-g_{j}({\mathbf{y}})\rvert$ $\leq G\left\|{\mathbf{x}}-{\mathbf{y}}\right\|$ (17)
@@
for all ${\boldsymbol{\xi}}$ and Mean Squared Lipschitz (MSL)
${\mathbb{E}}[\lvert f({\mathbf{x}},{\boldsymbol{\xi}})-f({\mathbf{y}},{\boldsymbol{\xi}})\rvert]$ $\leq G\left\|{\mathbf{x}}-{\mathbf{y}}\right\|$ (19)
for all ${\boldsymbol{\xi}}$.
@@
$\langle\nabla g_{j}({\tilde{\mathbf{x}}}),{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\rangle$ $\leq-\rho$ $\forall~{}j\in{\bar{\mathcal{J}}}$ :=$:=\{k\mid g_{k}({\tilde{\mathbf{x}}})\geq-\chi\}$ (20a)
$\langle\nabla h_{i}({\tilde{\mathbf{x}}}),{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\rangle$ $\leq-\rho$ $\forall~{}i\in{\bar{\mathcal{I}}}$ :=$:=\{l\mid h_{l}({\tilde{\mathbf{x}}})\geq-\chi\}.$ (20b)
@@ III-C2 Assumption on surrogate functions
Assumptions A8,A9, and A10 restrict the choice of the surrogate functions and have also been used in [21, 15]. Examples of surrogate functions satisfying Assumptions A8-A9 can be found in [47, 21] (see Sec III-D). An implication of Assumptions A3 and A9 is that the function $\phi_{j}({\mathbf{x}}):=g_{j}({\mathbf{x}})-{\tilde{g}}_{j}({\mathbf{x}},{\mathbf{x}}_{t})$ is $2L$-smooth so that
@@
$\langle\nabla h_{i}({\tilde{\mathbf{x}}}),{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\rangle\leq-\rho$ $\forall~{}~{}i\in{\bar{\mathcal{I}}}$ :=$:=\{i\mid h_{i}({\tilde{\mathbf{x}}})\geq-\chi\}.$
$\left\|{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\right\|_{2}\leq 1$
@@ III-D Examples of Surrogate Functions
For $L$-smooth constraint function $g_{j}$, a common choice of surrogate that satisfies all required conditions is the quadratic upper bound ${\tilde{g}}_{j}({\mathbf{x}},{\mathbf{z}})=g_{j}({\mathbf{z}})+\langle\nabla g_{j}({\mathbf{z}}),{\mathbf{x}}-{\mathbf{z}}\rangle+\frac{L}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}$. If we can write $g_{j}({\mathbf{x}})=g_{j}^{c}({\mathbf{x}})+g_{j}^{\bar{c}}({\mathbf{x}})$ where the first summand is a convex function and the second summand is an $L$-smooth non-convex one, we can use the surrogate ${\tilde{g}}_{j}({\mathbf{x}},{\mathbf{z}})=g_{j}^{c}({\mathbf{x}})+g_{j}({\mathbf{z}})+\langle\nabla g_{j}({\mathbf{z}}),{\mathbf{x}}-{\mathbf{z}}\rangle+\frac{L}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}$.
In addition to these two classes of surrogates, the function $f$ admits a few more structures. For instance, if $f({\mathbf{x}},{\boldsymbol{\xi}})=f_{1}(f_{2}({\mathbf{x}},{\boldsymbol{\xi}}))$ where $f_{1}$ is a convex function admits the surrogate ${\hat{f}}({\mathbf{x}},{\mathbf{y}},{\boldsymbol{\xi}})=f_{1}(f_{2}({\mathbf{y}},{\boldsymbol{\xi}})+\langle\nabla f_{2}({\mathbf{y}},{\boldsymbol{\xi}}),{\mathbf{x}}-{\mathbf{y}}\rangle)+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{y}}\right\|^{2}$. In the case of block convex functions, the variable ${\mathbf{x}}$ can be partitioned into disjoint blocks denoted by $\{{\mathbf{x}}_{k}\}_{k=1}^{K}$ and the objective function can be expressed as $f({\mathbf{x}},{\boldsymbol{\xi}})=\sum_{k=1}^{K}f^{k}({\mathbf{x}}_{k},{\mathbf{x}}_{-k},{\boldsymbol{\xi}})$ where each $f^{k}$ is convex in its first argument and ${\mathbf{x}}_{-k}$ collects all variables other than those in the block ${\mathbf{x}}_{k}$. For such functions, we can have the surrogate
${\hat{f}}({\mathbf{x}},{\mathbf{z}},{\boldsymbol{\xi}})=\sum_{k=1}^{K}f^{k}({\mathbf{x}}_{k},{\mathbf{z}}_{-k},{\boldsymbol{\xi}})+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}.$ (24)
@@ IV Complexity analysis
Lemma 1 provides a bound on gradient tracking error at the $(t+1)$-th iteration in terms of the tracking error at $t$-th iteration as well as the iterate progress norm. The proof follows from expanding the left-hand side of (31) and introducing $(1-\beta_{t+1})\nabla U({\mathbf{x}}_{t})$. Subsequently, using the smoothness of $f$, Assumption A5, and the definition of $G_{t+1}$, we obtain the desired bound. The full proof of this lemma is provided in Appendix A. Next we establish the descent lemma.
@@
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]\leq\frac{{\mathbb{E}}\left[\eta_{t}\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]}{2}-\frac{{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]}{4}$ (32)
Lemma 2 bounds the change in $F$ at every iteration. The proof of Lemma 2 starts with using the smoothness of $U$ and the convexity of $u$. Subsequent use of the optimality condition of $\mathcal{P}_{t}$ and Assumption A8 yields the desired bound. The full proof of this lemma is provided in Appendix B. Next, we show that the choice of the surrogate functions ${\tilde{g}}_{j}$ ensures that the iterates remain feasible for all $t\geq 1$.
@@
The iterates $\{{\mathbf{x}}_{t}\}_{t\geq 1}$ produced by Algorithm 1 are feasible for the original problem ($\mathcal{P}$)
@@
$\geq\frac{1}{{\mathbb{E}}\left[1/\eta_{T}\right]}{\mathbb{E}}\left[\sqrt{\sum_{t=1}^{T}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}}\right]^{2}\geq\frac{\bar{k}T\Delta_{T}^{2}}{(w+TG^{2})^{1/3}}.$ (42)
@@
$\leq{\bar{k}}^{3}\sum\nolimits_{t=1}^{T}\frac{G_{t+1}^{2}}{(G^{2}+\sum_{i=1}^{t+1}G_{i}^{2})}\leq{\bar{k}}^{3}\ln\Big{(}1+\sum\nolimits_{t=1}^{T+1}\frac{G_{t}^{2}}{G^{2}}\Big{)}\leq{\bar{k}}^{3}\ln(T+2)$ (43)
@@
$\Psi_{T+1}-\Psi_{1}$ $\leq-\frac{\bar{k}T\Delta_{T}^{2}}{8(w+TG^{2})^{1/3}}+\frac{c^{2}\bar{k}^{3}}{8L^{2}}\ln(T+2)$ (44)
@@
The bound in (47a) of Corollary 1 can be established along the lines of Theorem 1 by lower bounding the RHS of (41) and using $\eta_{T}~{}\geq~{}\tfrac{{\bar{k}}}{(w+TG^{2})^{1/3}}$. The bound in (47b) of Corollary 1 can be established on similar lines. Here, instead of dropping the first term on the right of (105), we drop the second term instead, and obtain a bound on the first term by proceeding along similar lines.
It is important to note that in Lemma 1, the last term of (31) affects the upper bounds, causing the convergence rate to depend on a logarithmic term. However, if the adaptive step sizes in (11) are replaced with fixed step sizes $\eta=\tfrac{{\bar{k}}}{T^{1/3}}$ and $\beta=c\eta^{2}$, the upper bound in Theorem 1 improves to $\mathcal{O}(T^{-2/3})$. Building on this observation, we present the following corollary, whose proof is omitted for brevity.
@@
Under Assumptions A1-A2, A4-A6 and A8, if $0<\bar{k}\leq 1$, $\tfrac{1}{4\bar{k}^{2}}\geq c>8L^{2}$, and $\mu>\tfrac{3}{4}$, with $\eta=\frac{{\bar{k}}}{T^{1/3}}$ and $\beta=c\eta^{2}$, it holds that
$\Delta_{T}\leq\sqrt{\frac{M^{\prime}_{T}}{T^{2/3}}}={\mathcal{O}}\left(\frac{1}{T^{1/3}}\right)$ (48a)
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]\leq\frac{M^{\prime}_{T}}{T^{2/3}}={\mathcal{O}}\left(\frac{1}{T^{2/3}}\right)$ (48b)
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|\right]\leq\sqrt{\frac{\bar{M^{\prime}}_{T}}{d^{\prime}T^{2/3}}}={\mathcal{O}}\left(\frac{1}{T^{1/3}}\right)$ (48c)
where $M^{\prime}_{T}:=8\frac{B_{1}}{{\bar{k}}}+\frac{\sigma^{2}}{2L^{2}\bar{k}^{2}}+\frac{c^{2}\bar{k}^{2}}{L^{2}}G^{2}$, and $d^{\prime}=\tfrac{1}{2L^{2}}(c-8L^{2})$
The proof of Corollary 2 follows on the lines of Theorem 1 and has been detailed in Appendix E.
@@
Lemma 4 implies that a strong version of Slater’s constraint qualification holds for ($\mathcal{P}_{t}$). In general, one cannot directly assume the existence of such as Slater point because the intermediate subproblems depend on ${\mathbf{x}}_{t}$ which is generated by the algorithm. The proof of Lemma 4 is provided in Appendix C and follows from the smoothness and Lipschitz continuity properties of ${\tilde{g}}_{j}$ and $h_{i}$.
@@
To prove Lemma 5 we use the convexity of ${\tilde{f}}$, ${\tilde{g}}_{j}$, $h_{i}$, and $u$, as well as the results in Lemma 4 and Assumption A10. The proof of Lemma 5 is provided in Appendix D. Finally, we state the main result, which is regarding the existence of an $\epsilon$-optimal KKT point.
@@
$\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})$ $=\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t}))$ (52)
$=\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{z}}_{t+1}$ (53)
@@
${\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}}$ $+\nabla{\mathbf{h}}({{\hat{{\mathbf{x}}}}_{t}}){\hat{{\boldsymbol{\nu}}}}_{t}=-\nabla{\tilde{{\mathbf{g}}}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t}){\hat{{\boldsymbol{\lambda}}}}_{t}-\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-{\mathbf{z}}_{t+1}$ (54)
@@
$=\nabla U({{\hat{{\mathbf{x}}}}_{t}})-{\mathbf{z}}_{t+1}+\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\Big{(}\nabla{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})-\nabla{\tilde{{\mathbf{g}}}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})\Big{)}{\hat{{\boldsymbol{\lambda}}}}_{t}=:{\boldsymbol{\pi}}_{t}$ (55)
@@
$\leq L\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|+\left\|{\boldsymbol{\varepsilon}}_{t}\right\|=L\left\|{\boldsymbol{\delta}}_{t}\right\|+\left\|{\boldsymbol{\varepsilon}}_{t}\right\|$ (56)
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\sum_{j=1}^{J}\left\|[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}(\nabla g_{j}({{\hat{{\mathbf{x}}}}_{t}})-\nabla{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t}))\right\|\leq 2L\left\|{\hat{{\boldsymbol{\lambda}}}}_{t}\right\|_{1}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|.$ (58)
@@
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}$ $\leq\tfrac{2}{T}\sum_{t=1}^{T}\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|\right]^{2}+\tfrac{2}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\varepsilon}}_{t}\right\|\right]^{2}$ (60)
$\leq\left(M_{T}\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}+\frac{\bar{M}_{T}}{d}\right)\frac{2(w+TG^{2})^{1/3}}{T}={\tilde{\mathcal{O}}}(T^{-2/3})$ (61)
@@
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})]$ $\geq-\frac{2B_{U}L^{2}}{\rho^{2}}\frac{M_{T}(w+TG^{2})^{1/3}}{T}={\tilde{\mathcal{O}}}(T^{-2/3}).$ (64)
@@
Under a non-adaptive step-size setting with $\eta=\frac{{\bar{k}}}{T^{1/3}}$ and $\beta=c\eta^{2}$, we state the following corollary, which follows along the lines of Theorem 2 and is detailed in Appendix E.

### v4

@@ III-C1 Assumptions on ($\mathcal{P}$)
$\left\|\nabla f({\mathbf{x}},{\boldsymbol{\xi}})-\nabla f({\mathbf{y}},{\boldsymbol{\xi}})\right\|$ $\leq L\left\|{\mathbf{x}}-{\mathbf{y}}\right\|,$ (12)
@@
$\lvert g_{j}({\mathbf{x}})-g_{j}({\mathbf{y}})\rvert$ $\leq G\left\|{\mathbf{x}}-{\mathbf{y}}\right\|.$ (17)
@@
and Mean Squared Lipschitz (MSL)
${\mathbb{E}}[\lvert f({\mathbf{x}},{\boldsymbol{\xi}})-f({\mathbf{y}},{\boldsymbol{\xi}})\rvert]$ $\leq G\left\|{\mathbf{x}}-{\mathbf{y}}\right\|.$ (19)
@@
$\langle\nabla g_{j}({\tilde{\mathbf{x}}}),{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\rangle$ $\leq-\rho$ $\forall~j\in{\bar{\mathcal{J}}}$ :=$:=\{k\mid g_{k}({\tilde{\mathbf{x}}})\geq-\chi\},$ (20a)
$\langle\nabla h_{i}({\tilde{\mathbf{x}}}),{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\rangle$ $\leq-\rho$ $\forall~i\in{\bar{\mathcal{I}}}$ :=$:=\{l\mid h_{l}({\tilde{\mathbf{x}}})\geq-\chi\},$ (20b)
@@ III-C2 Assumption on surrogate functions
[…]
@@
$\langle\nabla h_{i}({\tilde{\mathbf{x}}}),{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\rangle\leq-\rho$ $\forall~~i\in{\bar{\mathcal{I}}}$ :=$:=\{i\mid h_{i}({\tilde{\mathbf{x}}})\geq-\chi\}$
$\left\|{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\right\|_{2}\leq 1.$
@@ III-D Examples of Surrogate Functions
For $L$-smooth constraint function $g_{j}$, a common choice of surrogate that satisfies all required conditions is the quadratic upper bound ${\tilde{g}}_{j}({\mathbf{x}},{\mathbf{z}})=g_{j}({\mathbf{z}})+\langle\nabla g_{j}({\mathbf{z}}),{\mathbf{x}}-{\mathbf{z}}\rangle+\frac{L}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}$. If we can write $g_{j}({\mathbf{x}})=g_{j}^{c}({\mathbf{x}})+g_{j}^{\bar{c}}({\mathbf{x}})$ where the first summand is a convex function and the second summand is an $L$-smooth non-convex one, we can use the surrogate ${\tilde{g}}_{j}({\mathbf{x}},{\mathbf{z}})=g_{j}^{c}({\mathbf{x}})+g_{j}^{\bar{c}}({\mathbf{z}})+\langle\nabla g_{j}^{\bar{c}}({\mathbf{z}}),{\mathbf{x}}-{\mathbf{z}}\rangle+\frac{L}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}$.
In addition to these two classes of surrogates, the function $f$ admits a few more structures. For instance, if $f({\mathbf{x}},{\boldsymbol{\xi}})=f_{1}(f_{2}({\mathbf{x}},{\boldsymbol{\xi}}))$ where $f_{1}$ is a convex function admits the surrogate ${\hat{f}}({\mathbf{x}},{\mathbf{y}},{\boldsymbol{\xi}})=f_{1}(f_{2}({\mathbf{y}},{\boldsymbol{\xi}})+\langle\nabla f_{2}({\mathbf{y}},{\boldsymbol{\xi}}),{\mathbf{x}}-{\mathbf{y}}\rangle)+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{y}}\right\|^{2}$. In the case of block convex functions, the variable ${\mathbf{x}}$ can be partitioned into disjoint blocks denoted by $\{{\mathbf{x}}_{k}\}_{k=1}^{K}$ and the objective function can be expressed as $f({\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{k},{\boldsymbol{\xi}})$ where $f$ is convex in its first argument. For such functions, we can have the surrogate
${\hat{f}}({\mathbf{x}},{\mathbf{z}},{\boldsymbol{\xi}})=\sum_{k=1}^{K}\left(f({\mathbf{x}}_{k},{\mathbf{z}}_{-k},{\boldsymbol{\xi}})+\frac{\mu}{2}\left\|{\mathbf{x}}_{k}-{\mathbf{z}}_{k}\right\|^{2}\right),$ (24)
where ${\mathbf{z}}_{-k}$ collects all variables other than those in the block ${\mathbf{x}}_{k}$.
@@ IV Complexity analysis
Lemma 1 provides a bound on gradient tracking error at the $(t+1)$-th iteration in terms of the tracking error at $t$-th iteration as well as the iterate progress norm. The proof follows from expanding the left-hand side of (31) and introducing $(1-\beta_{t+1})\nabla U({\mathbf{x}}_{t})$. Subsequently, using the smoothness of $f$, Assumption A5, and the definition of $G_{t+1}$, we obtain the desired bound. The full proof of this lemma is provided in Appendix C. Next we establish the descent lemma.
@@
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]\leq\frac{{\mathbb{E}}\left[\eta_{t}\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]}{2}-\frac{{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]}{4}.$ (32)
Lemma 2 bounds the change in $F$ at every iteration. The proof of Lemma 2 starts with using the smoothness of $U$ and the convexity of $u$. Subsequent use of the optimality condition of $\mathcal{P}_{t}$ and Assumption A8 yields the desired bound. The full proof of this lemma is provided in Appendix D. Next, we show that the choice of the surrogate functions ${\tilde{g}}_{j}$ ensures that the iterates remain feasible for all $t\geq 1$.
@@
The iterates $\{{\mathbf{x}}_{t}\}_{t\geq 1}$ produced by Algorithm 1 are feasible for the original problem ($\mathcal{P}$).
@@
$\geq\frac{1}{{\mathbb{E}}\left[1/\eta_{T}\right]}{\mathbb{E}}\left[\sqrt{\sum_{t=1}^{T}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}}\right]^{2}\geq\frac{\bar{k}T\Delta_{T}^{2}}{(w+TG^{2})^{1/3}},$ (42)
@@
$\leq{\bar{k}}^{3}\sum\nolimits_{t=1}^{T}\frac{G_{t+1}^{2}}{(G^{2}+\sum_{i=1}^{t+1}G_{i}^{2})}\leq{\bar{k}}^{3}\ln\Big{(}1+\sum\nolimits_{t=1}^{T+1}\frac{G_{t}^{2}}{G^{2}}\Big{)}$
$\leq{\bar{k}}^{3}\ln(T+2).$ (43)
@@
$\Psi_{T+1}-\Psi_{1}$ $\leq-\frac{\bar{k}T\Delta_{T}^{2}}{8(w+TG^{2})^{1/3}}+\frac{c^{2}\bar{k}^{3}}{8L^{2}}\ln(T+2).$ (44)
@@
The bound in (112b) of Corollary 1 can be established along the lines of Theorem 1 by lower bounding the RHS of (41) and using $\eta_{T}~\geq~\tfrac{{\bar{k}}}{(w+TG^{2})^{1/3}}$. The bound in (112c) of Corollary 1 can be established on similar lines. Here, instead of dropping the first term on the right of (34), we drop the second term instead, and obtain a bound on the first term by proceeding along similar lines.
It is important to note that in Lemma 1, the last term of (31) affects the upper bounds, causing the convergence rate to depend on a logarithmic term. However, if the adaptive step sizes in (11) are replaced with non-adaptive choices, the bounds in Theorem 1 and Corollary 1 can be recovered under a milder assumption, as formalized in the following corollary. The detailed proof is provided in Appendix E.
@@
Using non-adaptive step size, namely $\eta_{t}=\tfrac{{\bar{k}}}{(w+t)^{1/3}}$ and $\beta_{t+1}=c\eta_{t}^{2}$, the bounds stated in Theorem 1 and Corollary 1 continue to hold under the weaker MSS Assumption A6, with slightly modified constant coefficients.
@@
Lemma 4 implies that a strong version of Slater’s constraint qualification holds for ($\mathcal{P}_{t}$). In general, one cannot directly assume the existence of such as Slater point because the intermediate subproblems depend on ${\mathbf{x}}_{t}$ which is generated by the algorithm. The proof of Lemma 4 is provided in Appendix A and follows from the smoothness and Lipschitz continuity properties of ${\tilde{g}}_{j}$ and $h_{i}$.
@@
To prove Lemma 5 we use the convexity of ${\tilde{f}}$, ${\tilde{g}}_{j}$, $h_{i}$, and $u$, as well as the results in Lemma 4 and Assumption A10. The proof of Lemma 5 is provided in Appendix B. Finally, we state the main result, which is regarding the existence of an $\epsilon$-optimal KKT point.
@@
$\nabla$ ${\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})$
$=\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t}))$ (51)
$=\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{z}}_{t+1}.$ (52)
@@
${\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}}$ $+\nabla{\mathbf{h}}({{\hat{{\mathbf{x}}}}_{t}}){\hat{{\boldsymbol{\nu}}}}_{t}=-\nabla{\tilde{{\mathbf{g}}}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t}){\hat{{\boldsymbol{\lambda}}}}_{t}-\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-{\mathbf{z}}_{t+1}.$ (53)
@@
$=\nabla U({{\hat{{\mathbf{x}}}}_{t}})-{\mathbf{z}}_{t+1}+\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$
$+\Big{(}\nabla{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})-\nabla{\tilde{{\mathbf{g}}}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})\Big{)}{\hat{{\boldsymbol{\lambda}}}}_{t}=:{\boldsymbol{\pi}}_{t}.$ (54)
@@
$\hskip-28.45274pt\leq L\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|+\left\|{\boldsymbol{\varepsilon}}_{t}\right\|=L\left\|{\boldsymbol{\delta}}_{t}\right\|+\left\|{\boldsymbol{\varepsilon}}_{t}\right\|.$ (55)
@@
$\leq\sum_{j=1}^{J}\left\|[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}(\nabla g_{j}({{\hat{{\mathbf{x}}}}_{t}})-\nabla{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t}))\right\|$
$\leq 2L\left\|{\hat{{\boldsymbol{\lambda}}}}_{t}\right\|_{1}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|.$ (57)
@@
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}$
$~~~~\leq\tfrac{2}{T}\sum_{t=1}^{T}\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|\right]^{2}+\tfrac{2}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\varepsilon}}_{t}\right\|\right]^{2}$ (59)
$~~~~\leq\left(M_{T}\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}+\frac{M_{T}}{d}\right)\frac{2(w+TG^{2})^{1/3}}{T}$
$~~~~={\tilde{\mathcal{O}}}(T^{-2/3}).$ (60)
@@
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})]$ $\geq-\frac{2B_{U}L^{2}}{\rho^{2}}\frac{M_{T}(w+TG^{2})^{1/3}}{T}$
$={\tilde{\mathcal{O}}}(T^{-2/3}).$ (63)
@@
Under a non-adaptive step-size setting with $\eta_{t}=\frac{{\bar{k}}}{(w+t)^{1/3}}$ and $\beta_{t+1}=c\eta_{t}^{2}$, we state the following corollary, which follows along the lines of Theorem 2 and whose proof is detailed in Appendix E.